OpenStudy (anonymous):

Undetermined Coefficients... y" + 10 y' + 16 y = 0

OpenStudy (anonymous):

roots = -2,-8

OpenStudy (anonymous):

Yp=0 ? Yh = c1 exp(-2t) + c2 exp(-8t) ?

OpenStudy (amistre64):

you have a homogenous equation so the homogenous is the solution

OpenStudy (anonymous):

now the test had initial conditions which I can't remember. I got to this point but couldn't remember how to apply the initial conditions. I remember y' = 4 lets say x = 0 and y =2 How do I apply the initial conditions

OpenStudy (amistre64):

take the derivatives of y as needed and solve the system that resutls y = a exp(-2t) + b exp(-8t) y' = -2a exp(-2t) - b exp(-8t) assuming you meant t=0 .... 2 = a exp(0) + b exp(0) 4 = -2a exp(0) - b exp(0)

OpenStudy (amistre64):

pfft, i deleted my 8

OpenStudy (anonymous):

Now its easy, thx. I frustrated with myself now, haha

OpenStudy (amistre64):

:) you did fine